# Weighted algebraic topology, II

### 05 May 2026 v1

@@ 3.3. Invariance and relative norms
(43) ${\underline{\mu}}(x)=\mu(0,x),\qquad\qquad\mu(x,y)={\underline{\mu}}(y-x).$
In fact, if $\mu\colon X\!\times\!X\to{\bf w}$ is an invariant $\mu$-metric, the associated mapping ${\underline{\mu}}\colon X\to{\bf w}$ satisfies (42)
$\begin{array}[]{l}{\underline{\mu}}(x)+{\underline{\mu}}(y)=\mu(0,x)+\mu(0,y)=\mu(0,x)+\mu(x,x+y)\\[3.0pt] \hskip 18.49988pt\hskip 18.49988pt\;\;\;\;\;\geqslant\;\mu(0,x+y)={\underline{\mu}}(x+y).\end{array}$
Conversely, assuming that ${\underline{\mu}}\colon X\to{\bf w}$ satisfies (42), the associated mapping $\mu\colon X\!\times\!X\to{\bf w}$ is obviously invariant up to translations and
$\mu(x,y)+\mu(y,z)={\underline{\mu}}(y-x)+{\underline{\mu}}(z-y)\,\geqslant\,{\underline{\mu}}(z-x)=\mu(x,z).$
(c) A product of abelian groups with $\mu$-metrics invariant up to translations is of the same kind. For instance all $\delta\mathbb{R}^{n}$, $\mu\mathbb{R}^{n}$, $\delta_{0}\mathbb{R}^{n}$ have a $\mu$-metric invariant up to translations.
(d) A $\mu$-metric $\mu$ on $X$ which is invariant up to translations and linear (as in $\mu\mathbb{R}$) amounts to a homomorphism of abelian groups
(44) $\varphi\colon X\to\mathbb{R},\qquad\qquad\varphi(x+y)=\varphi(x)+\varphi(y),$
which is at the same time its potential (null at $0_{X}$) and its relative norm:
(45) $\mu(x,y)=\varphi(y)-\varphi(x)=\varphi(y-x).$
In fact, if $\mu$ is invariant up to translations and linear, ${\underline{\mu}}(x)=\mu(0,x)=\varphi(x)$, and the relation (45) says that $\varphi$ is a homomorphism. Conversely, (45) trivially implies that $\varphi$ is a $\mu$-norm for $\mu$.
@@ 3.4. Discrete and chaotic metrics
(a) Every set $X$ can be equipped with the discrete $\mu$-metric $\mu_{d}$ (the greatest one), which is positive and symmetric
(46) $\mu_{d}(x,x)=0,\qquad\mu_{d}(x,y)=\infty\qquad\qquad(\mbox{for }x\neq y),$
@@
giving the discrete $\mu$-metric space $\mu_{d}X$, and the left adjoint to the forgetful functor $\mu\mathsf{Mtr}\to\mathsf{Set}.$ In $\delta\mathsf{Mtr}$ these issues will be written $\delta_{d}$ and $\delta_{d}X$. (Of course the associated topology is discrete.)
A discrete $\mu$-metric space is a sum of $\delta$-singletons. (For a sum of $\mu$-singletons, see 3.5(e).)
@@
(b) Every set $X$ has a chaotic $\mu$-metric and a chaotic $\delta$-metric (the least ones), which are symmetric
(47) $\mu_{c}(x,y)={\overline{\infty}},\qquad\delta_{c}(x,y)=0\qquad\qquad(x,y\in X).$
@@
The former gives the chaotic $\mu$-metric space $\mu_{c}X$, and the right adjoint to the forgetful functor $\mu\mathsf{Mtr}\to\mathsf{Set}$. The latter gives the chaotic $\delta$-metric space $\delta_{c}X$, and the right adjoint to $\delta\mathsf{Mtr}\to\mathsf{Set}$. (In both cases the associated topology is chaotic.)
@@
(c) We say that a $\mu$-space is affordable if $\mu(x,x^{\prime})<\infty$, for all pairs of points. The same term applies to $\delta$-spaces.
@@
An affordable $\mu$-space $X$ that has some points $z,z^{\prime}$ with $\mu(z,z^{\prime})={\overline{\infty}}$ is chaotic: $\mu(x,y)\leqslant\mu(x,z)+\mu(z,z^{\prime})+\mu(z^{\prime},y)={\overline{\infty}}$, for all $x,y$ (since all values are $\infty$).
3.5. Flat metrics
(a) Extending chaotic $\mu$-spaces, we come back to the flatness property, introduced in 2.2(b). A $\mu$-space $Z$ is flat if it satisfies the following equivalent properties
(i) $\mu(z,z)={\overline{\infty}}$, for all $z\in Z$,
(ii) $\mu(z,z^{\prime})=\pm\infty$, for all $z,z^{\prime}\in Z$.
Indeed, if (i) holds, $\mu(z,z^{\prime})\leqslant{\overline{\infty}}+\mu(z,z^{\prime})$, whence $\mu(z,z^{\prime})=\pm\infty$. The converse is also obvious, as the only possible values of $\mu(z,z)$ are 0 and $\overline{\infty}$}.
[…]
(48) $\mu(x,y)={\overline{\infty}}\;\mbox{ if }\>x\leqslant y,\qquad\mu(x,y)=\infty\;\text{ otherwise}.$
The metric is invariant up to translations (and associated to the $\mu$-norm ${\underline{\mu}}(x)={\overline{\infty}}$ if $x\geqslant 0$, and ${\underline{\mu}}(x)=\infty$ otherwise).
In fact, each preordered set $X$ has such a flat $\mu$-metric ${\rm mt}_{\infty}(X)$, see 4.3. If the preorder is an equivalence relation we get a symmetric flat $\mu$-metric, characterised in (d).
(c) A flat $\mu$-space $Z$ is a disjoint union $\raisebox{1.93747pt}{${\mbox{$\bigcup$}}$}\,Z_{i}$ of maximal chaotic $\mu$-spaces, the equivalence classes of the relation $z\!\sim\!z^{\prime}$ if $\mu(z,z^{\prime})={\overline{\infty}}=\mu(z^{\prime},z)$. Transitivity is obvious, from $\mu(z,z^{\prime})+\mu(z^{\prime},z^{\prime\prime})\geqslant\mu(z,z^{\prime\prime})$. Let us note that, if $z\nsim z^{\prime}$, then $\mu(z,z^{\prime})$ or $\mu(z^{\prime},z)$ is $\infty$.
(d) A symmetric flat $\mu$-space $Z$ has an even simpler structure: the previous disjoint union is a sum $\raisebox{1.93747pt}{${\mbox{$\sum$}}$}\,Z_{i}$ of indecomposable chaotic $\mu$-spaces ($z\nsim z^{\prime}$ implies $\mu(z,z^{\prime})=\mu(z^{\prime},z)=\infty$). Therefore, in $Z$:
[…]
(e) In particular, a sum $Z$ of $\mu$-singletons has a symmetric flat structure
(50) $\mu(z,z)={\overline{\infty}},\qquad\mu(z,z^{\prime})=\infty\qquad\qquad(\mbox{for }z\neq z^{\prime}).$
3.6. Proposition (Symmetric $\mu$-spaces).
(a) A symmetric $\mu$-space $X$ decomposes in the sum
(51) $\begin{array}[]{c}X=X_{0}+Z,\\[5.0pt] X_{0}=\{x\in X\,|\,\mu(x,x)=0\},\;\;\;\;Z=\{z\in X\,|\,\mu(z,z)={\overline{\infty}}\},\end{array}$
where $X_{0}$ is a symmetric $\delta$-metric space, called the positive component of $X$, and $Z$ is a symmetric flat $\mu$-space, the flat part of $X$. The latter can be further decomposed in a sum $\raisebox{1.93747pt}{${\mbox{$\sum$}}$}\,Z_{i}$ of indecomposable chaotic $\mu$-metric spaces, as in 3.5(d).
(b) A symmetric $\mu$-space $X$ which is affordable is positive or chaotic. (The empty space is both.)
(c) The non-affordable symmetric $\mu$-space $\delta\top+\mu\top$ is neither positive nor chaotic; it is already decomposed as above.
Proof.
Let $X$ be a symmetric $\mu$-space. First we note that, for $z,z^{\prime}\in X$
(52) $\mu(z,z^{\prime})<0\;\implies\;\mu(z,z)=\mu(z^{\prime},z^{\prime})=\mu(z,z^{\prime})={\overline{\infty}}.$
In fact, the assumption $\mu(z,z^{\prime})<0$ makes $\mu(z,z)=0$ impossible, in the symmetric case: $\mu(z,z)\leqslant\mu(z,z^{\prime})+\mu(z^{\prime},z)=2\mu(z,z^{\prime})<0.$
Therefore $\mu(z,z)=\mu(z^{\prime},z^{\prime})={\overline{\infty}}$; then the subspace $\{z,z^{\prime}\}$ is flat, and $\mu(z,z^{\prime})={\overline{\infty}}$.
(a) By (52), the points of $X_{0}$ have a positive $\mu$-metric. We have still to prove that, if $x\in X_{0}$ and $z\in Z$, then $\mu(x,z)=\infty$. In fact $\mu(x,z)\leqslant\mu(x,z)+\mu(z,z)$ can only be $\pm\infty$, but $\mu(x,z)<0$ is impossible, by (52).
(b) If there exist $z,z^{\prime}$ with $\mu(z,z^{\prime})<0$, we have $\mu(z,z)={\overline{\infty}}$, by (52), and we conclude with 3.4(c). ∎
@@ 4. Symmetrisation, topology and preorder
We show that the full subcategory $\mu\mathsf{Mtr}\subset\mu\mathsf{Mtr}$ of symmetric $\mu$-metric spaces (see 2.2(d)) is reflective and coreflective in $\mu\mathsf{Mtr}$, producing two associated ‘symmetric’ topologies. In the positive case, the reflector was used in [G3, G4] while the coreflector was used in [Lw]. Similarly, a $\mu$-metric space has a reflective and a coreflective preorder, defined in 4.3.
@@
The ‘indeterminate’ relation of relative metrics to topology, discussed in the Introduction, is also investigated in 4.8.
@@ 4.1. The reflective symmetric metric and topology
(a) The reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr}$ is based on the greatest symmetric $\mu$-metric $\hat{\mu}\leqslant\mu$ (already used in [G3, G4] in the positive case, and written $\delta$), which will be called here the reflective symmetric $\mu$-metric of $\mu$
(53) $\begin{array}[]{c}{}^{\,{\scriptstyle{\wedge}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr},\qquad(X,\mu)^{\,{\scriptstyle{\wedge}}\,}=(X,\hat{\mu}),\qquad\hat{\mu}=\mu{\,{\scriptstyle{\wedge}}\,}\mu^{{\rm op}},\\[5.0pt] \hat{\mu}(x,x^{\prime})={\rm inf}_{\underline{x}}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,(\mu(x_{j-1},x_{j}){\,{\scriptstyle{\wedge}}\,}\mu(x_{j},x_{j-1}))),\\[5.0pt] \hat{\mu}(x,x^{\prime})\leqslant\mu(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\mu(x^{\prime},x),\end{array}$
where ${\underline{x}}=(x_{0},...,x_{n})$ is any finite sequence in $X$ with $x_{0}=x$, $x_{n}=x^{\prime}$ (a step-path from $x$ to $x^{\prime}$). The unit is the comparison $X\to X^{\,{\scriptstyle{\wedge}}\,}$, the counit is the identity $Y^{\,{\scriptstyle{\wedge}}\,}=Y$, for a symmetric $\mu$-space. (The inclusion functor $\mu\mathsf{Mtr}\subset\mu\mathsf{Mtr}$ is left understood.)
Restricting to the positive case, we obtain the reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\delta\mathsf{Mtr}\to\;!\delta\mathsf{Mtr}$, based on the greatest symmetric $\delta$-metric $\delta^{\,{\scriptstyle{\wedge}}\,}\leqslant\delta$. One can replace $\mu\mathsf{Mtr}$ by $\mu_{\infty}\mathsf{Mtr}$, and $\delta\mathsf{Mtr}$ by $\delta_{\infty}\mathsf{Mtr}$.
Loosely speaking, (53) ‘adjusts’ the symmetric function $\mu(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\mu(x^{\prime},x)$ to satisfy the triangular inequality in $X$. However, if this function already satisfies it, the meet $\mu{\,{\scriptstyle{\wedge}}\,}\mu^{{\rm op}}$ can be computed pointwise on $X\!\times\!X$:
(54) $\hat{\mu}(x,x^{\prime})=\mu(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\mu(x^{\prime},x).$
(b) Composing with the functor ${\rm tp}\colon!\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ of (38) we have a functor $\rm tp$$^{\,{\scriptstyle{\wedge}}\,}$, also written $^{\,{\scriptstyle{\wedge}}\,}$
(55) $^{\,{\scriptstyle{\wedge}}\,}={\rm tp}^{\,{\scriptstyle{\wedge}}\,}\colon\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top}\qquad\qquad(^{\,{\scriptstyle{\wedge}}\,}\colon\mu\mathsf{Mtr}\to\mathsf{Top}).$
which equips a $\mu$-metric space with the reflective (symmetric) topology, associated to $\hat{\mu}$.
(c) For the basic examples of 3.1(a):
@@
- $(\mu\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\mu_{c}\mathbb{R}$ and $(\delta_{0}\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\delta_{c}\mathbb{R}$ have a chaotic metric and chaotic topology.
@@
The first case above (and its consequences on cartesian powers, tensor powers and substructures) shows that the reflector is appropriate for the goals of [G3, G4]. The second shows that we cannot rely on it to get a useful topology on $\mu\mathbb{R}$.
@@ 4.2. The coreflective symmetric metric and topology
(a) The coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr}$ is based on the least symmetric $\mu$-metric $\check{\mu}\geqslant\mu$, which will be called the coreflective symmetric $\mu$-metric of $\mu$
(56) $\begin{array}[]{ccc}{}^{\,{\scriptstyle{\vee}}\,}\colon\mu\mathsf{Mtr}\to\;!\mu\mathsf{Mtr},&&(X,\mu)^{\,{\scriptstyle{\vee}}\,}=(X,\check{\mu}),\\[5.0pt] \check{\mu}=\mu{\,{\scriptstyle{\vee}}\,}\mu^{{\rm op}},&&\check{\mu}(x,x^{\prime})=\mu(x,x^{\prime}){\,{\scriptstyle{\vee}}\,}\mu(x^{\prime},x).\end{array}$
@@
Let us note that $\check{\mu}$ is computed pointwise on $X\!\times\!X$ because $\mu$-metrics are closed under joins. Moreover:
(57) $\hat{\mu}\leqslant\mu\leqslant\check{\mu},\qquad\hat{\mu}\leqslant\mu^{{\rm op}}\leqslant\check{\mu}.$
Restricting to the positive case, we obtain the coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\delta\mathsf{Mtr}\to\;!\delta\mathsf{Mtr}$, based on the least symmetric $\delta$-metric $\delta^{\,{\scriptstyle{\vee}}\,}\geqslant\delta$. Again, one can replace $\mu\mathsf{Mtr}$ by $\mu_{\infty}\mathsf{Mtr}$, and $\delta\mathsf{Mtr}$ by $\delta_{\infty}\mathsf{Mtr}$.
@@
(b) Composing with the functor ${\rm tp}\colon!\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ of (38) we have a functor $\rm tp$$^{\,{\scriptstyle{\vee}}\,}$, also written $^{\,{\scriptstyle{\vee}}\,}$
(58) $^{\,{\scriptstyle{\vee}}\,}={\rm tp}^{\,{\scriptstyle{\vee}}\,}\colon\mu_{\infty}\mathsf{Mtr}\to\mathsf{Top}\qquad(^{\,{\scriptstyle{\vee}}\,}\colon\mu\mathsf{Mtr}\to\mathsf{Top}),$
which equips a $\mu$-metric space with the coreflective (symmetric) topology, associated to $\check{\mu}$. For a $\mu$-space $X$ we have the following ‘finer’ relations, for metrics and topologies
@@
- $\mu\mathbb{R}^{\,{\scriptstyle{\vee}}\,}$ and $\delta_{0}\mathbb{R}^{\,{\scriptstyle{\vee}}\,}$ have the euclidean metric $\check{\mu}(x,x^{\prime})=\delta_{0}^{\,{\scriptstyle{\vee}}\,}(x,x^{\prime})=|x-x^{\prime}|$ and the euclidean topology.
The coreflector is thus of interest for $\mu\mathbb{R}$, and of little use for $\delta\mathbb{R}$. We shall find convenient ways to use both the reflector and the coreflector (see 4.5); or none of them, as far as topology is concerned (in Part III).

### 13 May 2026 v2

@@ 3.3. Invariance and real norms
(43) ${\underline{\mu}}(x)=\rho(0,x),\qquad\qquad\rho(x,y)={\underline{\mu}}(y-x).$
In fact, if $\rho\colon X\!\times\!X\to{\bf w}$ is an invariant $\rho$-metric, the associated mapping ${\underline{\mu}}\colon X\to{\bf w}$ satisfies (42)
$\begin{array}[]{l}{\underline{\mu}}(x)+{\underline{\mu}}(y)=\rho(0,x)+\rho(0,y)=\rho(0,x)+\rho(x,x+y)\\[3.0pt] \hskip 18.49988pt\hskip 18.49988pt\;\;\;\;\;\geqslant\;\rho(0,x+y)={\underline{\mu}}(x+y).\end{array}$
Conversely, assuming that ${\underline{\mu}}\colon X\to{\bf w}$ satisfies (42), the associated mapping $\rho\colon X\!\times\!X\to{\bf w}$ is obviously invariant up to translations and
$\rho(x,y)+\rho(y,z)={\underline{\mu}}(y-x)+{\underline{\mu}}(z-y)\,\geqslant\,{\underline{\mu}}(z-x)=\rho(x,z).$
(c) A product of abelian groups with $\rho$-metrics invariant up to translations is of the same kind. For instance all $\delta\mathbb{R}^{n}$, $\rho\mathbb{R}^{n}$, $\delta_{0}\mathbb{R}^{n}$ have a $\rho$-metric invariant up to translations.
(d) A $\rho$-metric $\rho$ on $X$ which is invariant up to translations and linear (as in $\rho\mathbb{R}$) amounts to a homomorphism of abelian groups
(44) $\Phi\colon X\to\mathbb{R},\qquad\qquad\Phi(x+y)=\Phi(x)+\Phi(y),$
which is at the same time its potential (null at $0_{X}$) and its real norm
(45) $\rho(x,y)=\Phi(y)-\Phi(x)=\Phi(y-x).$
In fact, if $\rho$ is invariant up to translations and linear, ${\underline{\mu}}(x)=\rho(0,x)=\Phi(x)$, and the relation (45) says that $\Phi$ is a homomorphism. Conversely, (45) trivially implies that $\Phi$ is a $\rho$-norm for $\rho$.
@@ 3.4. Discrete and chaotic metrics
(a) Every set $X$ can be equipped with the discrete $\rho$-metric $\rho_{D}$ (the greatest one), which is positive and symmetric
(46) $\rho_{D}(x,x)=0,\qquad\rho_{D}(x,y)=\infty\qquad\qquad(\mbox{for }x\neq y),$
@@
giving the discrete $\rho$-metric space $\rho_{D}X$, and the left adjoint to the forgetful functor $\rho\mathsf{Mtr}\to\mathsf{Set}.$ In $\delta\mathsf{Mtr}$ these issues will be written $\delta_{D}$ and $\delta_{D}X$. (Of course the associated topology is discrete.)
A discrete $\rho$-metric space is a sum of $\delta$-singletons.
@@
(b) Every set $X$ has a chaotic $\rho$-metric and a chaotic $\delta$-metric (the least ones), which are symmetric
(47) $\rho_{C}(x,y)={\overline{\infty}},\qquad\delta_{C}(x,y)=0\qquad\qquad(x,y\in X).$
@@
The former gives the chaotic $\rho$-metric space $\rho_{C}X$, and the right adjoint to the forgetful functor $\rho\mathsf{Mtr}\to\mathsf{Set}$. The latter gives the chaotic $\delta$-metric space $\delta_{C}X$, and the right adjoint to $\delta\mathsf{Mtr}\to\mathsf{Set}$. (In both cases the associated topology is chaotic.)
@@
(c) We say that a $\rho$-space is affordable if $\rho(x,x^{\prime})<\infty$, for all pairs of points. The same term applies to $\delta$-spaces.
@@
An affordable $\rho$-space $X$ that has some points $z,z^{\prime}$ with $\rho(z,z^{\prime})={\overline{\infty}}$ is chaotic: $\rho(x,y)\leqslant\rho(x,z)+\rho(z,z^{\prime})+\rho(z^{\prime},y)={\overline{\infty}}$, for all $x,y$ (since all values are $\infty$).
@@ 4. Symmetrisation, topology and preorder
We show that the full subcategory $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$ of symmetric $\rho$-metric spaces (see 2.2(d)) is reflective and coreflective in $\rho\mathsf{Mtr}$, producing two associated ‘symmetric’ topologies. In the positive case, the reflector was used in [G3, G4] while the coreflector was used in [L1]. Similarly, a $\rho$-metric space has a reflective and a coreflective preorder, defined in 4.3.
@@
The ‘indeterminate’ relation of real valued metrics to topology, discussed in the Introduction, is also investigated in 4.8.
@@ 4.1. The reflective symmetric metric and topology
(a) The reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is based on the greatest symmetric $\rho$-metric $\hat{\rho}\leqslant\rho$ (already used in [G3, G4] in the positive case, and written $\delta$), which will be called here the reflective symmetric $\rho$-metric of $\rho$
(48) $\begin{array}[]{c}{}^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},\qquad(X,\rho)^{\,{\scriptstyle{\wedge}}\,}=(X,\hat{\rho}),\qquad\hat{\rho}=\rho{\,{\scriptstyle{\wedge}}\,}\rho^{{\rm op}},\\[5.0pt] \hat{\rho}(x,x^{\prime})={\rm inf}_{\underline{x}}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\,(\rho(x_{j-1},x_{j}){\,{\scriptstyle{\wedge}}\,}\rho(x_{j},x_{j-1}))),\\[5.0pt] \hat{\rho}(x,x^{\prime})\leqslant\rho(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(x^{\prime},x),\end{array}$
where ${\underline{x}}=(x_{0},...,x_{n})$ is any finite sequence in $X$ with $x_{0}=x$ and $x_{n}=x^{\prime}$ (a step-path from $x$ to $x^{\prime}$). The unit is the comparison $X\to X^{\,{\scriptstyle{\wedge}}\,}$, the counit is the identity $Y^{\,{\scriptstyle{\wedge}}\,}=Y$, for a symmetric $\rho$-space. (The inclusion functor $\rho\mathsf{Mtr}\subset\rho\mathsf{Mtr}$ is left understood.)
Restricting to the positive case, we obtain the reflector $^{\,{\scriptstyle{\wedge}}\,}\colon\delta\mathsf{Mtr}\to\;!\delta\mathsf{Mtr}$, based on the greatest symmetric $\delta$-metric $\delta^{\,{\scriptstyle{\wedge}}\,}\leqslant\delta$. One can replace $\rho\mathsf{Mtr}$ by $\rho_{\infty}\mathsf{Mtr}$, and $\delta\mathsf{Mtr}$ by $\delta_{\infty}\mathsf{Mtr}$.
Loosely speaking, (48) ‘adjusts’ the symmetric function $\rho(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(x^{\prime},x)$ to satisfy the triangular inequality in $X$. However, if this function already satisfies it, the meet $\rho{\,{\scriptstyle{\wedge}}\,}\rho^{{\rm op}}$ can be computed pointwise on $X\!\times\!X$
(49) $\hat{\rho}(x,x^{\prime})=\rho(x,x^{\prime}){\,{\scriptstyle{\wedge}}\,}\rho(x^{\prime},x).$
(b) Composing with the functor ${\rm tp}\colon!\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ of (38) we have a functor $\rm tp$$^{\,{\scriptstyle{\wedge}}\,}$, also written $^{\,{\scriptstyle{\wedge}}\,}$
(50) $^{\,{\scriptstyle{\wedge}}\,}={\rm tp}^{\,{\scriptstyle{\wedge}}\,}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}\qquad\qquad(^{\,{\scriptstyle{\wedge}}\,}\colon\rho\mathsf{Mtr}\to\mathsf{Top}).$
which equips a $\rho$-metric space with the reflective (symmetric) topology, associated to $\hat{\rho}$.
(c) For the basic examples of 3.1(a)
@@
- $(\rho\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\rho_{C}\mathbb{R}$ and $(\delta_{0}\mathbb{R})^{\,{\scriptstyle{\wedge}}\,}=\delta_{C}\mathbb{R}$ have a chaotic metric and chaotic topology.
@@
The first case above (and its consequences on cartesian powers, tensor powers and substructures) shows that the reflector is appropriate for the goals of [G3, G4]. The second shows that we cannot rely on it to get a useful topology on $\rho\mathbb{R}$.
@@ 4.2. The coreflective symmetric metric and topology
(a) The coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr}$ is based on the least symmetric $\rho$-metric $\check{\rho}\geqslant\rho$, which will be called the coreflective symmetric $\rho$-metric of $\rho$
(51) $\begin{array}[]{ccc}{}^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\;!\rho\mathsf{Mtr},&&(X,\rho)^{\,{\scriptstyle{\vee}}\,}=(X,\check{\rho}),\\[5.0pt] \check{\rho}=\rho{\,{\scriptstyle{\vee}}\,}\rho^{{\rm op}},&&\check{\rho}(x,x^{\prime})=\rho(x,x^{\prime}){\,{\scriptstyle{\vee}}\,}\rho(x^{\prime},x).\end{array}$
@@
Let us note that $\check{\rho}$ is computed pointwise on $X\!\times\!X$ because $\rho$-metrics are closed under joins. Moreover
(52) $\hat{\rho}\leqslant\rho\leqslant\check{\rho},\qquad\hat{\rho}\leqslant\rho^{{\rm op}}\leqslant\check{\rho}.$
Restricting to the positive case, we obtain the coreflector $^{\,{\scriptstyle{\vee}}\,}\colon\delta\mathsf{Mtr}\to\;!\delta\mathsf{Mtr}$, based on the least symmetric $\delta$-metric $\delta^{\,{\scriptstyle{\vee}}\,}\geqslant\delta$. Again, one can replace $\rho\mathsf{Mtr}$ by $\rho_{\infty}\mathsf{Mtr}$, and $\delta\mathsf{Mtr}$ by $\delta_{\infty}\mathsf{Mtr}$.
@@
(b) Composing with the functor ${\rm tp}\colon!\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}$ of (38) we have a functor $\rm tp$$^{\,{\scriptstyle{\vee}}\,}$, also written $^{\,{\scriptstyle{\vee}}\,}$
(53) $^{\,{\scriptstyle{\vee}}\,}={\rm tp}^{\,{\scriptstyle{\vee}}\,}\colon\rho_{\infty}\mathsf{Mtr}\to\mathsf{Top}\qquad(^{\,{\scriptstyle{\vee}}\,}\colon\rho\mathsf{Mtr}\to\mathsf{Top}),$
which equips a $\rho$-metric space with the coreflective (symmetric) topology, associated to $\check{\rho}$. For a $\rho$-space $X$ we have the following ‘finer’ relations, for metrics and topologies:
@@
- $\rho\mathbb{R}^{\,{\scriptstyle{\vee}}\,}$ and $\delta_{0}\mathbb{R}^{\,{\scriptstyle{\vee}}\,}$ have the euclidean metric $\check{\rho}(x,x^{\prime})=\delta_{0}^{\,{\scriptstyle{\vee}}\,}(x,x^{\prime})=|x-x^{\prime}|$ and the euclidean topology.
The coreflector is thus of interest for $\rho\mathbb{R}$, and of little use for $\delta\mathbb{R}$. We shall find convenient ways to use both the reflector and the coreflector (see 4.5); or none of them, as far as topology is concerned (in Part III).